famous productions performs London shows. the average show sells 1,300 tickets at $60 per ticket. There are 140 shows per year.

Famous Productions performs London shows. The companies. The average show has a cast of 65, each earing a net average of $340 per show. The cast is paid ather each show. The other variable cost is a program-prinieg coat of 58 por queet Areal fxod costs average show sells 1 300 kees at Seo per tcket There ษข 140 shows per yer No a dio ral shows can be held as the heater is aso used byother proucton total $728.000 Read the reguirements Requirement 1. Compute revenue and variable costs for each show Select the formula and enter the amounts to comoute sales revenue for each show Sales revenue per show Select the formula and enter the amounts to compute variable coets for each show Compuile the variable coets per show for oach cost separabely, and then compute the total variable costs per show Varable costs per show Cost of programs Cost of performers Total variable costs Requirement 2 Use the equation approach to compute the number of shows Famous Productions must perform each year to break even First, select the formuia to compute the nequired sales in units to break even Target proft Rearrange the formula you detemined above and oompute the required number of shows to break even The rumber of shows needed annualy to break even is Choose from any ist or enter any number in the input fields and then continue to the next question
Famous Protutons perbms Lenon shows. The average show sels 1,300 sants at $60 ticket. There are 140 show s per year No addnond shows can be held asthe theater saiso used by other production comparies. The average show has a cast of 65, each earning a net average of $340 per show The cast is paid afler each show The other variable cost is a program-printing cost of $8 per quest. Annual fxed costs total $T28,000 Road the requirements The number of shows needed annualy to break even is Requirement 3. Use the contribution margin ratio approach to compute the number of shows needed each year to eam a profit of $5,607,500Is this profe gol reaister Give your neasoning Begin by showing the formula and then entering the amounts to calculate the required saies dollars to eam a proft of $5.687.500 (Round the required saies in dollars to the nearest whoie dolar. Round amounts in the formula to two decimal places, XoX.ox Abbreviation used: CMcontribution margin.) Required sales in doliars Now use the information given and the required sales in dolars computed in the previous step to delermine the required number of shows needed each year to eam a profit of $5,687500. (Round your answer up to the nearest whole number) The number of shows needed annually to eam a proft of $5,687,500 is s this profit goal realsse Give your reasoning The profit goal of $5,687,500 is since Famous Producions currently performs 140 shows a year
Requirement 4. Prepare Famous Productionss contribution margin income statement for 140 shows performed in 2018. Report only two categories of costs: variable and fixed Famous Productions Contribution Margin Income Statement Year Ended December 31, 2018 Operating Income (Loss)
Answer #1
Requirement 1 Compuation of Revenue & Variable Costs for ech show.
No. of ticket * price per ticket                           = Sales Revenue per show
1300*60                                                                = 78000
Cost of programms Programme printing cost of each Guest*No. of tickets =1300*8 = 10400
Cost of Performers No. of cast*average earning of each cast =65*340 = 22100
Total Variable cost per show 32500
Requirement 2
Sales Revenue*No. of shows-Variable cost*No. of shows-Fixed cost=Target profit
To Compute the required number of shows to breakeven is
To fin the number of shows to breakeven, the target profit has to be 0
Then No. of shows*(Sales revene - variable cost) = fixed cost
No. of shows*(78000-32500)=728000
No. of shows to breakeven is =728000/(78000-32500) = 16
Requirement 3 To earn a profit of 5687500
(Fixed cost+Targer profit)/Contribution margin ratio   = Required sales in dollars
(728000+5687500)/58.3333333%               = 10998000
Contribution margin ratio =Contribution/Sales%
Contribution margin ratio =(78000-32500)/78000%
58.33333333
Divide the Dollar sales by Revenue per show to arrive at No, of shows to earn the profit of $5687500
=10998000/78000 = 141
Yes This profit goal is realistic since the Famous production is already performing 140 shows a year.
Requirement 4 Sales Revenue 10920000
Variable costs 4550000
Fixed costs 728000
Operating Income 5642000